Q2) 1 Point for part a, 1 point for part b; total 2 points A company has developed a new fishing line, which the company claiPlease write the full derivation of the answers. Thank you!
Answer #1

a) Given that the null hypothesis is true that is the mean is 15, the probability of type I error is computed here as:

\alpha = P(\bar X < 14.9)

Converting it to a standard normal variable, we have here:

\alpha = P(Z < \frac{14.9 - 15}{\frac{0.5}{\sqrt{50}}})

\alpha = P(Z < - 1.4142)

Getting it from the standard normal tables, we have here:

\alpha = P(Z < - 1.4142) = 0.0786

Therefore 0.0786 is the required proability of type I error here.

b) For a true mean of 14.8, the probability of type II error that is beta is computed here as:

= Probability that the sample mean is greater than 14.9 given that the true mean is 14.8.

\beta_{\mu = 14.8} = P(\bar X > 14.9)

Converting it to a standard normal variable we have here:

\beta_{\mu = 14.8} = P(Z > \frac{14.9 - 14.8}{\frac{0.5}{\sqrt{50}}})

\beta_{\mu = 14.8} = P(Z > 1.4142)

Getting it from the standard normal tables, we have here:

\beta_{\mu = 14.8} = 0.0786

Therefore 0.0786 is the required type II error probability here.

For a true mean value of 14.9, the probability of type II error is computed here as:

\beta_{\mu = 14.8} = P(Z > \frac{14.9 - 14.9}{\frac{0.5}{\sqrt{50}}})

\beta_{\mu = 14.8} = P(Z > 0) = 0.5

Therefore 0.5 is the required type II error probability here.